# Evaluating the impact of racial disproportionality in Stop & Search on expressive crimes in London

### v1

@@ 5 Discussion
In this paper, we introduce a Bayesian procedure for semi-parametric partially linear structural regression which allows us to place a specific prior to the parameter of interest. In addition, this approach also addresses issues in high-dimensional scenarios. By integrating double machine learning method and sample splitting procedure into our Bayesian framework, it preserves key asymptotic properties and consistent estimation on the parameter of interest. In particular, we showed that computations following this paradigm yield valid posterior inference in terms of coverage according to Monahan and Boos (1992). Our approach also accommodates flexible machine learning models for treatment and outcome, mitigating the impact of model misspecification. In our application, we estimate the effect of impact of ethnic disproportionality in Stop & Search on expressive crimes in London with an informative prior, and conclude that: Stop & Search practices in London are disproportionally distributed among London boroughs and disproportionality targeting the Black community is mitigated by considering the proportion of the Black population for expressive crimes. The ethnic composition of demographics plays a significant role in affecting the identified disproportionality, with lower DI alleviated by higher percentage of black population in target boroughs.
Bayesian methods hold significant interest in applied scientific causal research. These methods enable direct probability statements regarding treatment effectiveness and facilitate sensitivity assessments with different prior or expert inputs. There remains many scopes for future research to explore various possibilities in Bayesian semi-parametric inference in high-dimensional settings. For example, the proposed methodology can be widely applied in other causal settings when the traditional Bayesian set-up requires over-specifying the model condition. Moreover, this Bayesian method, coupled with machine learning methods, also presents an important direction for advancement to accommodate complex data structures, such as time-varying treatments, longitudinal data, or hierarchical frameworks in high-dimensional scenarios.

### v2

@@ 6 Discussion
In this paper, we introduce a Bayesian procedure for semi-parametric partially linear structural regression which allows us to place a specific prior to the parameter of interest. In addition, this approach also addresses issues in high-dimensional scenarios. By integrating double machine learning method and sample splitting procedure into our Bayesian framework, it preserves key asymptotic properties and consistent estimation on the parameter of interest. In particular, we showed that computations following this paradigm yield valid posterior inference in terms of coverage according to Monahan and Boos (1992). Our approach also accommodates flexible machine learning models for treatment and outcome, mitigating the impact of model misspecification. In our application, we estimate the effect of impact of ethnic disproportionality in stop and search for expressive crimes in London with an informative prior, and conclude that stop and search practices in London are disproportionally distributed among London boroughs and disproportionality involving the Black community is mitigated by considering the proportion of the Black population for expressive crimes. The ethnic composition of demographics plays a significant role in affecting the identified disproportionality, with lower DI alleviated by higher percentage of black population in target boroughs. In other words, our analysis reveals that police are more aggressively policing Black individuals in areas that are predominantly white.
The estimand considered in our application differs from previous work such as that of Zhao et al. (2022), in which it does not contrast outcomes from minority groups with outcomes from non-minority groups. Also, the outcome of interest in this analysis is the event of a stop interaction, as opposed to the use of police force following a stop interaction or detainment as in Knox et al. (2020) and Zhao et al. (2022). Our DI compares the intensity of stop interactions involving Black individuals across different boroughs of London. In this sense, we are not examining whether Black individuals are treated worse than those of other races, but rather we are studying how Black individuals are treated in the different boroughs to detect geographic inequality. It could be that racial bias is severe in terms of how minority groups are treated compared to non-minority groups, but that it is done equally across the city. A future analysis involving individual-level data comprising different ethnic groups could entail estimating the estimands discussed in Zhao et al. (2022) using the proposed Bayesian double machine learning methodology. Moreover, we should note that residual spatial autocorrelation was not accounted for in this application. An important future extension would involve incorporating a Gaussian random Markov field component to the model to account for spatial dependence.
Bayesian methods hold significant interest in applied scientific causal research. These methods enable direct probability statements regarding treatment effectiveness and facilitate sensitivity assessments with different prior or expert inputs. There remains many scopes for future research to explore various possibilities in Bayesian semi-parametric inference in high-dimensional settings. For example, the proposed methodology can be widely applied in other causal settings when the traditional Bayesian set-up requires over-specifying the model condition. Moreover, this Bayesian method, coupled with machine learning methods, also presents an important direction for advancement to accommodate complex data structures, such as time-varying treatments, longitudinal data, interference, or hierarchical frameworks in high-dimensional scenarios.
@@ Acknowledgments
Appendix A Additional details regarding the Stop & Search dataset
A.1 Demographic data on ethnic groups
Exploratory data analysis on Stop & Search subjects’ demographics, especially the ethnic compositions, has found disproportionally distributed stop and search events among the ethnic groups. White people took up on average 40% over time (ranged from 38% to 42%), Black people took up on average 38% during the observation period (ranged from 36% to 40%), Asian people took up on average 17% (ranged from 14% to 18%), then Others at 5% (ranged from 4% to 6%).
Figure 4: Stop & Search Subjects Ethnic Groups Proportions
A.2 Data aggregation procedures
The data have been sourced from the Metropolitan Police and City of London police force and published in the Single Online Home National Digital Team under Open Government Licence v3.0. They are publicly available from https://data.police.uk/docs/method/stops-at-location/. Before publishing, the location coordinates of for each stop are anonymized (detailed methods of anonymization from the publisher can be found at https://data.police.uk/about/#location-anonymisation) and the age of the person stopped has been adjusted to a corresponding age group (e.g. 18-24).
In this research, 2.6 million stop and search incidents have been compiled for the observation period from January 2019 to December 2023, for seven legislative grounds (three of which are listed in Table 5) for 11 types of Searched Objects. Upon the aggregation of the data into three listed categories, drug objects related stop and search incidents took up 64% over the observation period, as compared to 12% for acquisitive objects and 24% for expressive objects. The searched subjects’ ethnic information have been collected from “officer defined ethnic group”.
Table 5: Legislative bases underlying the aggregation of stop and search interactions
Category Searched Objects Legislation